I'm trying to solve when a man knows he has lost 12.4 percent of his weight, and now weighs 249.6. What was his starting weight?
I was thinking something like
12.4 249.6
_____ = _____
100 x
Can't recall how to do this.

Answer by mananth(16946)

let weight be x
now 249.6
...
if it was 100
it would increase to 112.4
100/112.4 = x/249.6
original weight = x=249.6*100/112.4
x= 222.06